# Magnon Blockade with Skyrmion Qubit-Magnon Coupling in a Hybrid Quantum System

### v1

@@ III.2 Numerical Simulation of the Second-Order Correlation Function
Figure 5: The dependence of the second-order correlation function $g^{(2)}(0)$ on both the skyrmion driving field $\Omega_{s}/\gamma$ and the magnon driving field $\Omega_{d}/\gamma$. In (a), $g^{(2)}(0)$ is plotted as a function of $\Omega_{s}/\gamma$ for detuning values $\kappa/\gamma=0.15$, $\kappa/\gamma=0.3$, and $\kappa/\gamma=0.45$, $\Omega_{d}/\gamma=0.01$. In (b), the behavior of $g^{(2)}(0)$ is depicted as a function of $\Omega_{d}/\gamma$ under the same detuning conditions, $\Omega_{s}/\gamma=0.06$, while the other parameters are the same as those in Fig. 4.
In Fig. 5, the relationships among $\Omega_{d}$, $g_{ms}$, $\Omega_{s}$, and $\kappa$ are more clearly demonstrated. It shows that the system exhibits the strongest antibunching effect around $\Omega_{s}/\gamma\approx 0.06$ and $\Omega_{d}/\gamma\approx 0.01$. A lower decay rate ($\kappa/\gamma=0.15$) helps the system maintain a stronger antibunching effect, as indicated by the lower value of $g^{(2)}(0)$. As the decay rate increases, the antibunching effect weakens, but even at higher dissipation, $\log_{10}g^{(2)}(0)$ still reaches a good magnitude.
Figure 6: The equal-time second-order correlation function $g^{(2)}(0)$ is plotted as a function of the effective skyrmion qubit-magnon coupling strength $g_{ks}/\gamma$ under different thermal noise occupation numbers $n_{th}=(0,10^{-7},10^{-8})$. The parameters used in this calculation are $(\Delta,\Omega_{s},\Omega_{d})/\gamma=(10,0.06,0.01)$, while the other parameters are the same as those in Fig. 4.
Finally considering a bosonic environment at nonzero temperature, we can analyze the impact of ambient thermal noise on the magnon blockade. Fig. 6 illustrates the significant dependence of the second-order correlation function $g^{(2)}(0)$ on the coupling strength $g_{ms}/\gamma$, exhibiting different behaviors under various thermal magnon occupation numbers $n_{th}$. By numerically solving the master equation, we plotted the variation of $g^{(2)}(0)$ as a function of the qubit-magnon coupling strength $g_{ms}/\gamma$ for different thermal magnon occupation numbers $n=(0,10^{-8},10^{-7})$. With $g_{ms}=19.6$, the results show that the thermal magnon number significantly affects the magnon blockade effect. When $n_{th}=0$ (red solid line), the second-order correlation function reaches its minimum value, around $10^{-6}$, near $g_{ms}/\gamma\approx 20$, indicating that the system is in a strong blockade regime, where magnon excitations are significantly suppressed. For $n_{th}=10^{-8}$, $n_{th}=10^{-7}$, and $n_{th}=10^{-6}$ (green dashed line, blue dashed line, and purple dashed line, respectively), the magnon blockade effect is disrupted as the thermal magnon occupation number increases. Although antibunching behavior can still be observed near $g_{ms}/\gamma\approx 20$, the minimum value increases compared to the zero thermal occupation case, indicating that the presence of thermal magnons weakens the blockade effect. Thus, for a more pronounced observation of the magnon blockade, the YIG material must be cooled to low temperatures. The thermal magnon occupation number is given by $n_{\text{th}}=\left[\exp\left(\frac{\hbar\omega_{m}}{k_{B}T}\right)-1\right]^{-1}$, where $\omega_{m}=9.8$ GHz. At temperatures of 4.06mK, 4.64mK, the corresponding thermal occupation numbers are $10^{-8}$, $10^{-7}$, respectively. This figure demonstrates that adjusting both the thermal magnon occupation number $n_{th}$ and the coupling strength effectively controls the magnon blockade effect, influencing multi-particle excitations in the magnon mode, which is crucial for the design of quantum information processing and quantum control systems.
@@
Figure 7: The variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ is shown as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $g_{ms}/\gamma$. The probe field strength of the magnons is fixed at $\Omega_{d}/\gamma=0.01$, while other system parameters are set to $\kappa_{m}/\gamma=\kappa_{q}/\gamma=1$.
In the skyrmion-qubit system, information is stored in the quantum degree of helicity, and the logical states can be controlled by electric and magnetic fields, as reflected in $K_{0}$ and $E_{z}$, respectivelyPsaroudaki and Panagopoulos (2021).
For the case discussed above, where $E_{z}\approx 9.8\,\text{GHz}$ and $K_{0}\approx 0$, the computed $g^{(2)}(0)$ as a function of detuning, derived from numerical simulations under various driving strengths, demonstrates excellent agreement with the corresponding analytical predictions. This concordance substantiates the validity of the approximations employed in the theoretical framework. Furthermore, the absence of instability effects arising from small dissipation provides additional confirmation of the robustness of the adopted approximation.
In addition, for the scenario where $E_{z}\approx 0$ and $K_{0}\approx 9.8\,\text{GHz}$, the system is configured to exclusively drive the magnon mode, while the external driving on the skyrmion qubit is removed. The skyrmion is instead modulated via external magnetic and electric fields, with the modulation explicitly affecting the parameters $K_{0}$ and $E_{z}$. The Hamiltonian is written as
@@
Figure 8: The steady-state $\log_{10}g^{(2)}(0)$ is presented as a function of (a) the skyrmion-qubit coupling strength $g_{ms}$ and (b) the control field strength $\Omega_{s}/\gamma$, along with the dissipation rate $\kappa$, under the condition $\Delta/\gamma\approx\frac{1}{2}g_{ms}/\gamma$. In panel (a), $\Omega_{d}/\gamma=0.01$ is set, while in panel (b), $g_{ms}/\gamma=50.1$ is chosen, while the other parameters are the same as those in Fig. 7.
The coupling term $\tilde{g}_{ms}$ induces transitions between energy levels, with the coupling strength ranging from 50.1 MHz to 200 MHz, whereas the coupling term $g_{ms}$ causes a small shift in the energy levels of the skyrmion qubit, with the shift ranging from 19.6 MHz to 36 MHz Pan et al. (2024a). This shift is significantly smaller than the resonance frequencies of the hybrid system ($K_{0}\approx\omega_{m}\approx 9.8\,\text{GHz}$), and therefore, this term can be neglected. By employing the rotating-wave approximation (RWA), the Hamiltonian of the hybrid quantum system can be simplified. The effective Hamiltonian can then be written as
$\hat{H}_{\text{eff}}^{\prime}=$ $\,\Delta_{m}\hat{m}^{\dagger}\hat{m}+\frac{\Delta_{s}}{2}\tilde{\sigma}_{z}+\frac{g_{ms}}{2}(\hat{m}\tilde{\sigma}_{+}+\hat{m}^{\dagger}\tilde{\sigma}_{-})$ (14)
$+\Omega_{d}(\hat{m}^{\dagger}+\hat{m}),$
where $\Delta_{m}\equiv\omega_{m}-\omega_{d}$ ($\Delta_{s}\equiv K_{0}-\omega_{d}$) represents the detuning between the magnon mode and the driving field frequency, while $\Delta_{s}$ represents the detuning between the skyrmion qubit frequency and the driving frequency. Simultaneously, $\Delta=\Delta_{s}=\Delta_{m}$. The master equationis expressed as followsCarmichael (1999)
$\frac{\partial\rho}{\partial t}=$ $-i[H_{\text{eff}},\rho]+\frac{\kappa_{m}}{2}\mathcal{L}_{m}[\rho]+\frac{\kappa_{s}}{2}\mathcal{L}_{\sigma}[\rho],$ (15)
@@
The Fig. 7 shows the variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $g_{ms}/\gamma$. It can also be observed that under the optimal conditions where $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$, $\log_{10}g^{(2)}(0)$ reaches -2, indicating the occurrence of a strong magnon blockade effect. This results in a symmetric relationship, which differs slightly from the case where two drives are applied simultaneously.
@@
In the analysis of Fig. 8(a), we investigate the effects of the dissipation rate $k$ and the Skyrmion qubit-magnon coupling strength $g_{\text{ms}}$ on the system’s behavior. The results show that, for higher coupling strengths and lower dissipation rates $k$, the logarithmic value of the second-order correlation function $\log_{10}g^{(2)}(0)$ remains at a relatively low value (around -5), consistent with the two driving influences acting on the system. In Fig. 8(b), we further examine the impact of the driving strength $\Omega_{d}$ on the system. When the dissipation rate $k$ is fixed, small variations in the driving strength (e.g., $0.001\leq\Omega_{d}\leq 0.01$) do not significantly affect the system, and the logarithmic value of the second-order correlation function $\log_{10}g^{(2)}(0)$ remains stable. However, when the driving strength undergoes a large variation, the system exhibits pronounced fluctuations, and the second-order correlation function shows significant changes, with a marked enhancement of the magnon blockade effect. Notably, even when no external drive is applied to the Skyrmion qubit, the application of a drive solely to the magnons effectively induces a magnon blockade effect.
V Conclusion
We propose a scheme to realize a controlled magnon blockade effect in a hybrid skyrmion-magnon system. By numerically solving the quantum master equation describing the system’s Hamiltonian, we found the statistical properties of the magnon mode, particularly observing a pronounced magnon blockade effect through the second-order correlation function $g^{(2)}(0)$. We consider two scenarios: one where both the magnon and the skyrmion are driven, with $E_{z}\approx 9.8\,\text{GHz}$ and $K_{0}\approx 0$, and another where only the magnon is driven, with $E_{z}\approx 0$ and $K_{0}\approx 9.8\,\text{GHz}$.
In the first scenario, where both the magnon and skyrmion are driven and the skyrmion is modulated by a magnetic field, we can effectively control $g^{(2)}(0)$ by adjusting the coupling strength $g_{\text{ms}}$ and the driving fields $\Omega_{s}$ or $\Omega_{d}$, leading to the observation of quantum antibunching phenomena. The model supports the realization of magnon blockade across a wide range of coupling strengths and external driving fields. The coupling strength between the skyrmion qubit and the magnon mode is directly influenced by the vertical distance between the skyrmion and the surface and center of the YIG sphere, with a maximum coupling strength of up to 40 MHz Pan et al. (2024a). Under these conditions, when the thermal magnon occupation number is $n_{\text{th}}=0$, the second-order correlation function $g^{(2)}(0)$ can be reduced to the order of $10^{-7}$.
In the second scenario, where only the magnon is driven and the skyrmion is modulated by a magnetic field, the coupling strength $g_{\text{ms}}$ can reach up to 200 MHz Pan et al. (2024a). By appropriately tuning the coupling strength $g_{\text{ms}}$ and the driving field $\Omega_{d}$, the magnon blockade effect is also observed. Further analysis of the system’s dissipation, driving fields shows that, under suitable parameter settings, the magnon blockade effect is still observed, with $g^{(2)}(0)$ reaching approximately $10^{-2}$. We believe that with further optimization and experimental validation, this scheme will provide a theoretical foundation for the potential application of single-magnon emitters.
@@ Acknowledgments
We thank Feng-Yang Zhang, Yexiong Zeng, and Ye-Ting Yan for constructive discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12274053).

### v2

@@ III.3 Numerical Simulation of the Second-Order Correlation Function
Figure 6: The dependence of the second-order correlation function $g^{(2)}(0)$ on both the skyrmion driving field $\Omega_{s}/\gamma$ and the magnon driving field $\Omega_{d}/\gamma$. In (a), $g^{(2)}(0)$ is plotted as a function of $\Omega_{s}/\gamma$ for detuning values $\kappa/\gamma=0.15,0.30,0.45,1$, and $\Omega_{d}/\gamma=0.01$. In (b), the behavior of $g^{(2)}(0)$ is depicted as a function of $\Omega_{d}/\gamma$ under the same detuning conditions, $\Omega_{s}/\gamma=0.06$, while the other parameters are the same as those in Fig. 5.
In Fig. 6, we illustrate the relationship between $\Omega_{d}$, $g_{ms}$, $\Omega_{s}$, and $\kappa$, showing that the system achieves the strongest antibunching effect near $\Omega_{s}/\gamma\approx 0.06$ and $\Omega_{d}/\gamma\approx 0.01$. Stronger antibunching effects are observed at lower dissipation rates, such as $\kappa/\gamma=0.15$, which is evidenced by the significantly reduced values of $g^{(2)}(0)$. As the dissipation rate increases, the antibunching effect weakens; however, $g^{(2)}(0)$ remains at a sufficiently low value even under higher dissipation conditions. For instance, when $\kappa/\gamma=1.00$, $g^{(2)}(0)$ can reach $10^{-4}$.
Our approach to the MB effect aims to drive the magnon system into a quantum state where the first excited level is highly populated, while higher excited levels remain largely unoccupied, thereby suppressing multi-excitation events. This mechanism minimizes the second-order correlation function $g^{(2)}(0)$, a key indicator of the MB effect. It is expressed as $g^{(2)}(0)\approx 2P_{2}/P_{1}^{2}$, which aligns with Eqs. (8) and (17). Fig. 7 shows the population dynamics of the lowest Fock states of the magnon mode under optimal conditions. The population of the single-excitation level $P_{1}$ is approximately $10^{-2}$, while the population of the double-excitation state $P_{2}$ is around $10^{-11}$. With the direct coupling between magnons and qubits, a nearly perfect magnon blockade can be achieved Jin and Jing (2024), characterized by $P_{1}\sim 0.40$ and $g^{(2)}(0)\sim 10^{-5}$. These results provide strong evidence of MB and highlight its potential for precise quantum control in hybrid quantum systems.
Figure 7: The population dynamics of the lowest few Fock states of the magnon mode under the optimal conditions $\Delta/\gamma\approx\pm\frac{1}{2}g_{\text{ms}}/\gamma$ and $\Omega_{s}=6\Omega_{d}$. The parameters are $g_{\text{ms}}/\gamma=19.6$, $\Omega_{d}/\gamma=0.01$, and $\kappa/\gamma=0.15$. The initial state is $\rho(0)=|g^{\prime}0\rangle\langle g^{\prime}0|$.
@@
In the previous analysis, we investigated the case where the skyrmion qubit is solely modulated by an electric field ($E_{z}\approx 9.8\,\text{GHz}$ and $K_{0}\approx 0$) and observed a clear MB effect. This analysis provided valuable insights into the interaction mechanisms under such conditions. Moving forward, we consider another scenario where $E_{z}\approx 0$ and $K_{0}\approx 9.8\,\text{GHz}$. In this configuration, the system is designed to exclusively probe the magnon mode, with the external driving on the skyrmion qubit removed. The Hamiltonian is written as
@@
Following the approach used in the previous analysis for handling the Hamiltonian, The effective Hamiltonian can then be written as
$\hat{H}_{\text{eff}}^{\prime}=$ $\Delta_{m}\hat{m}^{\dagger}\hat{m}+\frac{\Delta_{s}}{2}\sigma_{z}+\frac{\tilde{g}_{ms}}{2}\left(\hat{m}\sigma_{+}+\hat{m}^{\dagger}\sigma_{-}\right)$
$+\Omega_{d}\left(\hat{m}^{\dagger}+\hat{m}\right),$
where $\Delta_{m}\equiv\omega_{m}-\omega_{d}$ ($\Delta_{s}\equiv K_{0}-\omega_{d}$) represents the detuning between the magnon mode and the driving field frequency, while $\Delta_{s}$ represents the detuning between the skyrmion qubit frequency and the driving frequency. Simultaneously, $\Delta=\Delta_{s}=\Delta_{m}$. The statistical properties of the magnon are investigated by numerically solving the master equationCarmichael (1999)
$\frac{\partial\rho}{\partial t}=$ $-i[\hat{H}_{\text{eff}}^{\prime},\rho]+\frac{\kappa_{m}}{2}\mathcal{L}_{m}[\rho]+\frac{\kappa_{s}}{2}\mathcal{L}_{\sigma}[\rho],$ (20)
@@
Figure 8: The variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ is shown as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $\tilde{g}_{ms}/\gamma$. The probe field intensity of the magnons is fixed at $\Omega_{d}/\gamma=0.01$, while other system parameters are set to $\kappa_{m}/\gamma=\kappa_{q}/\gamma=1$.
@@
The Fig. 8 shows the variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $g_{ms}/\gamma$. It can also be observed that under the optimal conditions where $\Delta/\gamma\approx\pm\frac{1}{2}\tilde{g}_{ms}/\gamma$, $\log_{10}g^{(2)}(0)$ reaches -2, indicating the occurrence of a strong MB effect.
Figure 9: The steady-state $\log_{10}g^{(2)}(0)$ is presented as a function of (a) the skyrmion-qubit coupling strength $g_{ms}$ and (b) the control field intensity $\Omega_{s}/\gamma$, along with the dissipation rate $\kappa$, under the condition $\Delta/\gamma\approx\pm\frac{1}{2}\tilde{g}_{ms}/\gamma$. In panel (a), $\Omega_{d}/\gamma=0.01$ is set, while in panel (b), $g_{ms}/\gamma=50.1$ is chosen, and $n_{th}=0$.
@@
In the analysis of Fig. 9(a), the effects of the dissipation rate $k$ and the skyrmion qubit-magnon coupling strength $\tilde{g}_{ms}$ on the system’s behavior are explored. The result indicates that, under high coupling strength and low dissipation rate $k$, $\log_{10}g^{(2)}(0)$ remains at a low value of approximately -5. In Fig. 9(b), the influence of the driving intensity $\Omega_{d}$ on the system is further investigated. When the dissipation rate $k$ is fixed, small variations in the driving intensity (e.g., $0.001\leq\Omega_{d}/\gamma\leq 0.01$) have little effect on the system, with $\log_{10}g^{(2)}(0)$ remaining stable at around -3.5. However, with a larger variation in the driving intensity, the system exhibits pronounced fluctuations, leading to noticeable changes in the second-order correlation function. For instance, as $\Omega_{d}/\gamma$ increases from 0.1 to 0.6, $\log_{10}g^{(2)}(0)$ rises from -3.5 to -1.5. Notably, even in the absence of an external drive applied to the skyrmion qubit, applying a drive solely to the magnons effectively induces a MB effect.
V conclusions
In conclusion, we propose a scheme to realize a high-purity single-magnon state in a skyrmion qubit-magnon system. We consider two scenarios: one where the magnon is probed and the skyrmion qubit is driven, with the skyrmion qubit modulated by an electric field ($E_{z}\approx 9.8\,\text{GHz}$ and $K_{0}\approx 0$); and another where only the magnon is probed, with the skyrmion qubit modulated by a magnetic field ($E_{z}\approx 0$ and $K_{0}\approx 9.8\,\text{GHz}$). In the first scenario, we derive the analytical expression for the second-order correlation function and the optimal conditions through theoretical analysis and numerical simulations. Specifically, the detuning is one-half of the transverse coupling strength of the skyrmion qubit, and the driving field strength applied to the skyrmion qubit is six times the probing field strength applied to the magnon. Under these optimal conditions, and with the thermal magnon occupation number $n_{\text{th}}=0$, a strong MB effect is observed, with the second-order correlation function $g^{(2)}(0)\sim 10^{-7}$. This demonstrates that detuning, driving field strength, and probing field strength significantly influence the enhancement of the MB effect. Moreover, we analyze the physical mechanism responsible for the magnon blockade and confirm the coexistence of CMB and UMB. In the second scenario, by appropriately tuning the coupling strength $g_{\text{ms}}$ and the probing field $\Omega_{d}$, the MB effect is also observed. Further analysis of the system’s dissipation, driving fields shows that, under suitable parameter settings, the MB effect is still observed, with $g^{(2)}(0)$ reaching approximately $10^{-2}$. These findings establish a theoretical framework for studying and advancing the potential realization of single-magnon emitters in hybrid quantum systems.
Additionally, the versatility of this platform allows for further exploration of its applicability in broader quantum technologies. By leveraging the tunability of coupling parameters and system configurations, this scheme may inspire the design of scalable quantum architectures and novel quantum devices. These advancements could contribute to enriching the functionality of hybrid quantum systems and unlocking new possibilities in magnon-based quantum technologies.
@@ Acknowledgments
We thank Feng-Yang Zhang, Mingsong Ding, and Ye-Ting Yan for constructive discussions. This work was supported by the National Natural Science Foundation of China (Grants No. 12274053).